A rectangle's length is twice its width. The perimeter is 180 ft. Find the area, and length and width.
Answer by TimothyLamb(4379)

x = length
y = width
---
x = 2y
2x + 2y = 180
---
put the system of linear equations into standard form
---
x - 2y = 0
2x + 2y = 180
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 60 ft
y = 30 ft
---
answer:
he will ride 2.25 hours on his out-bound leg ...
z = 24x
z = 24*2.25
z = 54
he will ride 54 miles downwind, at which point he must turn around and head back home
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---